OpenStudy (anonymous):

I will give a medal I really need help Find the distance between these points. A (5, 8), B(-3, 4) AB =

OpenStudy (anonymous):

what is the fuction?

jigglypuff314 (jigglypuff314):

\[distance formula = \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given the points (x1, y1) and (x2, y2) so when given the points (5, 8) and (-3, 4) plug in x1=5 , y1=8 , x2=-3 , y2=4

OpenStudy (anonymous):

Is that the answer ?

OpenStudy (anonymous):

you should apply the distance for yourself not ask for the answer.

jigglypuff314 (jigglypuff314):

you would plug in x1=5 , y1=8 , x2=-3 , y2=4 into the distance formula and simplify to get your answer :)

jigglypuff314 (jigglypuff314):

so \[\sqrt{(-3 - (5))^{2} + (4 - 8)^{2}}\] then simplify

OpenStudy (anonymous):

so 3-5=5 and 4-8=5 so then 5x5=25 so is the answer 25 ??? I feel like its wrong ? im not sure

jigglypuff314 (jigglypuff314):

... it is a -3 - 5 so like -(3+5) = and 8 - 4 does not equal 5 ...

OpenStudy (anonymous):

so 8x 8=64

OpenStudy (calculusxy):

Are there any answer choices?

jigglypuff314 (jigglypuff314):

-(3+5) = -8 (8-4) = 4 \[\sqrt{(-8)*(-8) + (4)*(4)} = \sqrt{(64) + (16)}\]

OpenStudy (anonymous):

no but i have to write the number in a sqaure root forme

jigglypuff314 (jigglypuff314):

then what is 64+16 then put that number under a squareroot

OpenStudy (anonymous):

that sqare root of 80 is 8.9

jigglypuff314 (jigglypuff314):

yes, but you said that they wanted it in square root form so you would probably just keep it as squareroot of 80

OpenStudy (anonymous):